OpenStudy (anonymous):

find the value of m such that the function y=e^mx is asolution of y`+2y=0

OpenStudy (kainui):

Do you know what y' is?

OpenStudy (mathmale):

If so, perform that operation on y=e^(mx).

OpenStudy (anonymous):

y` is the first dirvative for y with respect to x i perform that operation on y but i cant continue e^mx*(m+2)=0

OpenStudy (kainui):

If you have \(a*b=0\) then either \(a=0\) or \(b=0\)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle y=e^{mx} }\) \(\color{#000000}{ \displaystyle y'=me^{mx} }\) (via chain rule) \(\color{#000000}{ \displaystyle y'+ny=0\quad \longrightarrow \quad me^{mx} +n(e^{mx} )=0\quad \longrightarrow \quad (m+n)e^{mx} =0 }\) Using the Zero Property (if I recall the terminology correctly), \(\color{#000000}{ \displaystyle 0=e^{mx} }\) or \(\color{#000000}{ \displaystyle 0=m+n }\) Then you know that: \(\color{#000000}{ \displaystyle e^{mx}\ne 0 ~~\forall\{m,x\} }\), and \(\color{#000000}{ \displaystyle m+n=0\quad \Longrightarrow \quad \color{red}{m=-n} }\)